OpenStudy (loser66):

Please, check my stuff: determine the solution set to the system Ax =0 for the given matrix A

OpenStudy (loser66):

\[\left[\begin{matrix}1-i&2i\\1+i&-2\end{matrix}\right]\] Multiply row 1 by -i , I get \[\left[\begin{matrix}-1-i&2\\1+i&-2\end{matrix}\right]\] add row1+row2.

OpenStudy (loser66):

so, (1+i )x= 2y the solution set will be S= { (1+i)/2 t, t| t in R} but it's wrong. I don't know why??

OpenStudy (anonymous):

\[\left[\begin{matrix}1-i & 2i \\ 1+i & 2\end{matrix}\right]\left(\begin{matrix}a \\ b\end{matrix}\right)=\left(\begin{matrix}0 \\ 0\end{matrix}\right)\] multiply out and solve for a & b

OpenStudy (anonymous):

(1-i)(-i)=-i +i^2 = -1-i

OpenStudy (anonymous):

check if det[A] = 0 because if it's not 0 then the 0 vector is only solution.

OpenStudy (loser66):

so, what 's wrong with mine?

OpenStudy (loser66):

nope, it's not that, friend.

OpenStudy (anonymous):

? it's not what?

OpenStudy (anonymous):

yeah, det[A] = 0 so there's a non-trivial solution.

OpenStudy (loser66):

since the matrix has one 0 row, therefore, there is one free variable, the solution set is infinite set . the result from book said that it is S ={ (1-i)t,t|t is in R}, not as what I did

OpenStudy (anonymous):

take a row... (1+i)x = 2y solve for y. y = (1+i)/2 * x

OpenStudy (anonymous):

let x = t, then y = [(1+i)/2]t

OpenStudy (loser66):

the same with mine, but it's not right. so. that's why I want you check what's wrong.

OpenStudy (anonymous):

it should be the same thing... to check, just solve for x in terms of y... y=(1+i)/2 * x => (1+i)x=2y => x = 2y/(1+i) now to divide mult by conjugate... \[ x=\frac{2y\cdot(1-i)}{1+1}=(1-i)y\] now let t = y then x = (1-i)t

OpenStudy (anonymous):

make sense?

OpenStudy (loser66):

give me time to read, does my requirement make sense?

OpenStudy (anonymous):

yur requirement? not following

OpenStudy (loser66):

just joke. you just post and then asked me "make sense?" so I required time to read. hihihi

OpenStudy (anonymous):

yeah, sorry... please take your time and let me know.

OpenStudy (loser66):

ok, it makes sense to me now. thank you.

OpenStudy (anonymous):

you're welcome... i forget that i've already worked it out and of course it makes sense to me and then I forget you've not had the time to look and digest. sorry!

OpenStudy (loser66):

hehehe... that's my fault. thanks again

OpenStudy (anonymous):

no... it's mine!